# Deep Learning-based Intrusion Detection Systems: A Survey

### v1

@@ 4.1.3. Cloud Platform Tools
Cloud-native scenarios introduce different challenges compared with Windows or Linux platforms. For example, there are many different types of components such as containers, microservices, and Kubernetes (K8S) clusters, each of which generates its own logs with widely varying formats and contents. Additionally, components are basically characterized by dynamic expansion and contraction, making it hard to capture the whole data provenance.
To address the above challenges, Chen et al. (Chen et al., 2023) design a cloud log collection architecture on the basis of K8S, which is a central platform based on cloud-native technology. Josef et al. (Horalek et al., 2022) propose a log collection and analysis tool operated as Software as a Service (SaaS) in the cloud environment in Kubernetes technology, aiming to provide comprehensive logs across all microservices.
saBPF (Lim et al., 2021) is an extension tool of eBPF, aiming to deploy fully-configurable, high-fidelity, system-level audit mechanisms at the granularity of containers. saBPF is further developed with proof-of-concept intrusion detection system and access control mechanism to demonstrate its practicability.
Note that some collection tools in Windows and Linux platforms such as auditd (Grubb, 2008), sysdig (Kili, [n. d.]), and Tracee (Tracee, 2022) can also be applied in cloud computing environment. To distinguish, we only introduce collection tools tailored to the cloud platform here.
Cross-platform Tools
To effectively detect intrusions, an intuitive idea is to incorporate log data from various platforms to obtain a global view of the running system. However, it is impossible to merge these data directly due to the difference in their log formats.
DTrace (Gregg and Mauro, 2011) is a real-time dynamic tracing framework for troubleshooting kernel and application problems on production systems. It supports many platforms, including Linux, Windows, Solaris, macOS, FreeBSD, NetBSD, etc.
Support for Provenance Auditing in Distributed Environments (SPADE) (Gehani and Tariq, 2012) develops a novel provenance kernel that mediates between the producers and consumers of provenance information, and handles the persistent storage of records. Generally, it supports heterogeneous aggregating for system-level data provenance for data analysis across multiple platforms.
@@ 4.2. Log Storage
After collecting logs from raw data, the subsequent step is to store these logs (Zhang et al., 2023; Ding et al., 2023a). Data storage involves integration, management, maintenance, and retrieval of the data, mainly focusing on metrics such as reliability and efficiency. The storage cannot tamper with the completeness of data, enabling developers to precisely reconstruct history scenarios. Additionally, when utilizing the collected data, developers should pursue low latency and high throughput for efficient data queries.
Here, we will introduce two essential components for the data storage process: log storage systems and compression algorithms for these systems.
@@ 4.2.1. Log Storage Systems
The two most commonly used log storage systems are ELK (Ahmed et al., 2020) and Loki (Bautista et al., 2022). ELK is a powerful log management solution consisting of three open-source software components: Elasticsearch (Elastic, 2010), Logstash (Elastic, 2009), and Kibana (Elastic, 2013). Elasticsearch is the leading distributed, RESTful search and analytics data engine designed with speed and scalability. In Elasticsearch, everything is indexed with finite state transducers and thus it supports rapid searches across vast repositories. Meanwhile, Elasticsearch implements a powerful Domain Specific Language, making it convenient and flexible to retrieve data. Logstash is a server-side data preprocessing pipeline to collect and integrate data from multiple sources. It performs filtration and transformation on its data and then sends the data to a favorable “stash” (usually Elasticsearch for ELK). Kibana is a data analytics and visualization platform at both speed and scale. It provides straightforward user interfaces to create a variety of charts and dashboards to visualize the trends, distribution, or other information of the data stored in Elasticsearch. ELK is powerful enough to be applied in enterprise scenarios, however, its performance comes at a price. ELK sacrifices ease of configuration and installation, and may simultaneously introduce severe runtime overhead for its hosts.
In contrast, Loki is a lightweight logging system with low resource overhead developed by Grafana Labs. It is designed with simple operations and efficient storage. Instead of indexing everything of data like ELK does, Loki mainly creates indices grounded in log labels. Moreover, Loki is well suited for open-source monitoring and visualization tools such as Prometheus (Prometheus, 2014) and Grafana (Labs, 2014). Integrating these two tools enables Loki to construct a complete monitoring and log analysis platform for information systems.
With the development of cloud computing, log storage systems are usually deployed on the cloud and reinforced by blockchain methods. Cao et al. (Cao et al., 2021) propose a cloud-native log database named LogStore. It uses highly-scalable, cost-effective cloud object storage to overcome the bandwidth limitations that arise when writing large amounts of logs. In their work, an optimized column index structure called LogBlock has been designed to support full-text search queries. Huang (Huang, 2019) proposes a log storage framework that uses the Ethereum blockchain to store the hash of log files and employs smart contracts to create indices for log files. Xu et al. (Xu et al., 2023b) propose a log data storage scheme that combines blockchain networks and IPFS technology. IPFS is used to store a large amount of file log data and the blockchain system is combined to achieve decentralized and secure log storage. Pourmajidi et al. (Pourmajidi and Miranskyy, 2018) propose Logchain, a blockchain-based logging system. It encrypts and seals logs and adds them to a hierarchical ledger to prevent log tampering, providing an immutable platform for log storage.
@@ 4.2.2. Log Compression Algorithms
Logs are generated quickly and require significant memory usage. For example, it is measured that a browser can produce about 10 GB of log data each day (Ding et al., 2023a). Such oversize data thus should be compressed before storage. Typically, log compression algorithms can be categorized into two types: general-purpose algorithms and those specifically adapted to log data.
@@
Dictionary-based LZ77 in gzip (Foundation, 1992), LZMA in 7zip_lzma (Pavlov, 2001), LZSS in quickLZ (QuickLZ, 2006)
Sorting-based BWT in bzip2 (Seward, 1996), ST in szip (Schindler, 1999)
Statistical-based PPMD in 7zip_ppmd, DMC in ocamyd (Schwellinger, 2008)
General compression algorithms refer to algorithms to reduce the size of data while preserving the information of the data. It is worth noting that general compression algorithms are also capable of compressing data other than logs, namely, log data is only one type of data they can compress. General compression algorithms can be classified into three categories based on their principles (Yao et al., 2020):
@@
Table 2 presents representative algorithms of the above three types. Due to the indeterminacy of statistical techniques, this type of compression algorithm may introduce losses in compression. Yet the other two types of algorithms are generally lossless. By evaluating the compression performance on 9 log files and 2 natural language files, a study shows that some general compression algorithms can achieve high compression ratios for log data, and log data is even easier to compress than natural language data (Yao et al., 2020).
@@
Different from natural language data, log data usually has specific structures and formal expressions. Therefore, compression algorithms should be adjusted for log data to improve the compression performance further. Yao et al. (Yao et al., 2021) propose LogBlock, which obtains small log blocks before compression and then uses a generic compressor to compress logs. Liu et al. (Liu et al., 2019) propose Logzip, which employs clustering algorithms to iteratively extract templates from raw logs and then obtain coherent intermediate representations for compressing logs. Facing the massive data streams generated by cloud services, Lin et al. (Lin et al., 2015) propose Cowic, which utilizes the characteristics of logs for column compression and can independently analyze one or several columns in log entries. Wei et al. (Wei et al., 2021b) propose a parser-based log compression tool, LogReducer, which shows a high compression ratio and fast compression speed in large-scale log scenarios such as the Alibaba Cloud environment. They further propose LogGrep (Wei et al., 2023) for fine-grained organization of log data using static and runtime patterns. Rodrigues et al. (Rodrigues et al., 2021) propose the lossless compression tool CLP, aiming to quickly retrieve log data while meeting compression requirements. CLP proposes to combine domain-specific compression and search with a generic lightweight compression algorithm. Li et al. (Li et al., 2024d) conduct empirical research on log data and propose LogShrink to overcome their observed limitations by leveraging the commonality and variability of log data. LogBlock (Yao et al., 2021) is designed to help existing jobs perform better. It reduces duplicate logs by preprocessing log headers and rearranging log contents, thereby improving the compression ratio of log files. LogReducer (Yu et al., 2023) is a framework that combines log hotspot identification and online dynamic log filtering. Its non-intrusive design significantly reduces log storage and runtime overhead. $\mu$Slope (Wang et al., 2024) is a compression and search method for semi-structured log data. It achieves efficient storage and query performance through data segmentation, pattern extraction, and index-free design. Denum (Yu et al., 2024) significantly improves log compression rates by optimizing the compression of digital tokens in logs. It is an efficient log compression tool suitable for scenarios where you need to save storage space or transmission bandwidth.
@@ 4.3. Log Parsing
To analyze log data, DL-IDS should first parse the data into structured and standardized formats for unified processing. This is because log data often originates from multiple different devices such as terminals, sensors, and network devices, which will easily lead to differences in their expressions.
The data parsing focuses on application logs and is usually executed by data classification and template extraction. Data classification is to classify log data into several groups through data structures such as vector representations of the data. Each group constitutes a template, which is used to extract features from log data and construct the structured logs. Basically, the existing log parsers mainly have three categories: clustering-based parsers, pattern-based parsers, and heuristic-based parsers. Figure 5 depicts the taxonomy of log parsers.
@@
Clustering-based parsers classify data using clustering algorithms. Xiao et al. (Xiao et al., 2023) propose LPV, which employs a hierarchical clustering algorithm to incrementally group logs based on Euclidean distance. Pokharel et al. (Pokharel et al., 2023) convert log messages into bi-grams and directly cluster them using the DBSCAN algorithm. Hamooni et al. (Hamooni et al., 2016) present a rapid log pattern recognition approach named LogMine. It is implemented in the map-reduce framework for distributed platforms to process millions of log messages in seconds. LogCluster (Lin et al., 2016) reduces the number of logs that need to be manually checked and improves the accuracy of problem identification through log clustering and the use of knowledge bases. LogOHC (Yang et al., 2019) achieves efficient log template extraction by combining text vectorization and online hierarchical clustering technology. It has high accuracy and real-time performance and is particularly suitable for processing multi-source heterogeneous log data. METING (Coustié et al., 2020) provides a robust and efficient log parsing method through frequent n-gram mining and flexible log grouping strategy, which can effectively process various types of log data. Kimura et al. (Kimura et al., 2019) propose an active fault detection system based on log generation patterns. Active fault detection of network logs was achieved by automatically extracting log templates, constructing log feature vectors, aggregating features, and using machine learning classifiers.
@@
Frequency-based parsers discover patterns that exceed the frequency threshold and employ the mined patterns to parse logs. Sedki et al. (Sedki et al., 2022) propose the log parsing tool ULP, which combines string matching and local frequency analysis to efficiently parse large log files. This method first uses text processing methods to group log events, and then conducts frequency analysis on the log instances in each group to identify fixed and variable content in the logs. Dai et al. (Dai et al., 2020) propose Logram, which utilizes an n-gram dictionary for log parsing. For n-grams with a frequency below the threshold, Logram recursively converts to (n-1)-grams until a list of uncommon 2-grams is obtained. To mitigate the parameter sensitivity issue in log parsers, Dai et al. (Dai et al., 2023) further proposed an entropy-based log parser PILAR, which balances parsing accuracy and efficiency. Zhou et al. (Zhou and Su, 2023) propose the log parser Polo, which extracts log templates from logs using a prefix forest composed of a ternary search tree. Xu et al. (Xu et al., 2023a) propose a hybrid log parsing model called Hue, which performs parsing through user-adaptive methods. Prefix-Graph (Chu et al., 2021) is an efficient, adaptive, and universal log parsing method that can stably extract log templates without relying on domain knowledge and manual parameter tuning. FT-tree (Zhang et al., 2017) dynamically builds a template tree by identifying frequently appearing word combinations in system logs, supporting incremental learning. It is used to extract templates from switch system logs in data center networks.
@@
Heuristic-based parsers rely on empirical knowledge to classify log data. He et al. (He et al., 2017) propose the online log parsing method Drain, which employs a depth-fixed parsing tree to group the original logs and encodes them using specially designed parsing rules. Each new log is matched layer by layer and assigned to the group with the highest similarity at the leaf node. Le et al. (Le and Zhang, 2023) propose to use a hint-based few-sample learning algorithm, LogPPT, to capture log template patterns. Utilizing new prompt tuning methods and an adaptive random sampling algorithm, LogPPT performs well on multiple public datasets. Liu et al. (Liu et al., 2022b) propose the UniParser parser to address the issue of difficult processing of heterogeneous logs, using the Token Encoder and Context Encoder modules to learn log context features. Spell (Du and Li, 2016) is an efficient streaming log parsing method that can dynamically extract log patterns in online processing and significantly improve processing efficiency through pre-filtering steps. Logan (Agrawal et al., 2019) achieves efficient and scalable log parsing through distributed processing, LCS matching, dynamic matching tolerance, and periodic merging. USTEP (Vervaet et al., 2021) is an online log parsing method based on an evolutionary tree structure that can discover and encode new parsing rules. It achieves constant parsing time and can efficiently parse raw log messages in a streaming manner. MoLFI (Messaoudi et al., 2018) transforms the log message format recognition problem into a multi-objective optimization problem and uses the NSGA-II algorithm to automatically generate high-quality log message templates.
@@ 5. Intrusion Detection
The intrusion detection stage uncovers intrusions using the semantic-level information. This section classifies and summarizes the mainstream graph summarization (Section 5.1), attack detection (Section 5.2), and attack investigation (Section 5.3).
@@ 5.1. Graph Summarization
It is illustrated that stealthy malware will inevitably interact with the underlying OS and be captured by provenance monitoring systems (Wang et al., 2020), which is the reason why PIDS (a form of DL-IDS) work and flourish recently. Log data generated from provenance monitoring system is referred to as data provenance as mentioned. Offering advantages in high precision, however, DL-IDS using data provenance sacrifices memory performance. They are required to process a vast volume of log data, since data provenance records all trails of events from their creations to their current states at a very fine-grained level, even some of which is trivial.
Therefore, unlike network traffic and application logs used by the other forms of DL-IDS, data provenance is fine-grained, detailed, and rich in semantics. The aforementioned log storage systems (Section 4.2) and log compression algorithms (Section 4.2.2) are insufficient in improving the space efficiency of data provenance due to the absence of semantic information. They compress log data at the token level or the byte level, aiming to find duplicates that can be merged or re-represented.
To this end, graph summarization is widely investigated to incorporate semantic-level information to further reduce the size of log data. In graph summarization, data provenance is transformed into a provenance graph, of which the causal relations are utilized to build the semantic understanding of system activities. Referring to the definition of data provenance (Definition 3.7), the definition of provenance graph can be defined as follows:
@@
Based on provenance graphs, graph summarization approaches are allowed to reduce the size of log data by directly removing irrelevant events, aggregating similar events, gathering similar execution entities, etc. Due to the exploitation of specific causal relations (or semantic information), graph summarization approaches are capable of performing erasion confidently and straightforwardly. This categorizes them as a type of lossy reduction, whereas the aforementioned log storage systems and log compression algorithms are usually lossless. We note that some surveys (e.g., (Zipperle et al., 2022; Inam et al., 2023)) may interchangeably use graph summarization and log compression to identify the approaches that reduce the size of log data. In this work, we explicitly distinguish them and refer to the lossless reduction as compression and the opposite one as summarization, respectively.
@@
Table 3 presents representative data summarization approaches. In terms of their execution modes, we classify them into two categories: offline graph summarization and online graph summarization.

### v2

@@ 4.1.3. Cloud Platform Tools
Although some collection tools in Windows and Linux platforms such as auditd (Grubb, 2008), sysdig (Kili, [n. d.]), and Tracee (Tracee, 2022) can be applied in cloud computing environment, cloud-native scenarios introduce different challenges compared with Windows or Linux platforms. First, there are many different types of components such as containers, microservices, and Kubernetes (K8S) clusters in cloud platforms, each of which generates its own logs with varying formats and contents. Additionally, components are basically characterized by dynamic expansion and contraction, making it hard to capture complete log data. To address them, Chen et al. (Chen et al., 2023) design a cloud log collection architecture on the basis of K8S, which is a central platform based on cloud-native technology. Josef et al. (Horalek et al., 2022) propose a log collection and analysis tool operated as Software as a Service (SaaS) in the cloud environment in K8S technology, aiming to provide comprehensive logs across all microservices. saBPF (Lim et al., 2021) is an extension tool of eBPF, aiming to deploy fully-configurable, high-fidelity, system-level audit mechanisms at the granularity of containers. saBPF is further developed with proof-of-concept IDS and access control mechanism to demonstrate its practicability. ISDC (Mirnajafizadeh et al., 2024) is designed to eliminate the bottleneck between network infrastructure (where data is generated) and security application servers (where data is consumed), which prioritizes specific flows to effectively optimize resource consumption.
4.1.4. Cross-platform Tools
To effectively detect intrusions, an intuitive idea is to incorporate log data from various platforms to obtain a global view of the running system. DTrace (Gregg and Mauro, 2011) is a real-time dynamic tracing framework for troubleshooting kernel and application problems on production systems. It supports many platforms, including Linux, Windows, Solaris, macOS, FreeBSD, NetBSD, etc. Support for Provenance Auditing in Distributed Environments (SPADE) (Gehani and Tariq, 2012) develops a novel provenance kernel that mediates between the producers and consumers of provenance information, and handles the persistent storage of records. It supports heterogeneous aggregating for system-level data provenance for data analysis across multiple platforms.
@@ 4.2. Log Storage
The subsequent step of log collection is to store these logs (Ding et al., 2023a; Altinisik et al., 2023). We will introduce two essential components for data storage: log storage systems and compression algorithms for these systems.
@@ 4.2.1. Log Storage Systems
The two most commonly used log storage systems are ELK (Ahmed et al., 2020) and Loki (Bautista et al., 2022). ELK is a powerful log management solution consisting of three open-source software components: Elasticsearch (Elastic, 2010), Logstash (Elastic, 2009), and Kibana (Elastic, 2013). Elasticsearch (Elastic, 2010) is the leading distributed, RESTful search and analytics data engine designed with speed and scalability. Logstash (Elastic, 2009) is a server-side data preprocessing pipeline to collect and integrate data from multiple sources. Kibana (Elastic, 2013) is a data analytics and visualization platform at both speed and scale. ELK is powerful enough to be applied in enterprise scenarios, however, its performance comes at a price. ELK sacrifices ease of configuration and installation, and may simultaneously introduce severe runtime overhead for its hosts. In contrast, Loki (Bautista et al., 2022) is a lightweight logging system with low resource overhead developed by Grafana Labs. It is designed with simple operations and efficient storage. Instead of indexing everything of data like ELK does, Loki mainly creates indices grounded in log labels. Moreover, Loki is well suited for open-source monitoring and visualization tools such as Prometheus (Prometheus, 2014) and Grafana (Labs, 2014). Integrating these two tools enables Loki to construct a complete monitoring and log analysis platform for information systems.
@@ 4.2.2. Log Compression Algorithms
Logs are generated quickly and require significant memory usage. For example, it is measured that a browser can produce about 10 GB of log data each day (Ding et al., 2023a). Such oversize data should be compressed before storage. Log compression algorithms can be categorized into two types: general-purpose algorithms and those specifically adapted to log data.
@@
Dictionary-based LZ77 in gzip (Foundation, 1992), LZMA in 7zip_lzma (Pavlov, 2001), and LZSS in quickLZ (QuickLZ, 2006)
Sorting-based BWT in bzip2 (Seward, 1996) andST in szip (Schindler, 1999)
Statistical-based PPMD in 7zip_ppmd and DMC in ocamyd (Schwellinger, 2008)
General compression algorithms refer to algorithms to reduce the size of data (e.g., log data) by handling token-level or byte-level duplicates in the data. General compression algorithms can be classified into three categories based on their principles (Yao et al., 2020):
@@
Table 2 presents representative algorithms of the above three types. Due to the indeterminacy of statistical techniques, statistical-based compression algorithms may introduce losses in compression. Yet the other two types of algorithms are generally lossless. By validating 9 log files and 2 natural language files, a study (Yao et al., 2020) shows that some general compression algorithms can achieve high compression ratios for log data and log data is even easier to compress than natural language data.
@@
Different from natural language data, log data usually has specific structures and formal expressions that help further compression. Yao et al. (Yao et al., 2021) propose LogBlock, which obtains small log blocks before compression and then uses a generic compressor to compress logs. Liu et al. (Liu et al., 2019) propose Logzip, which employs clustering algorithms to iteratively extract templates from raw logs and then obtain coherent intermediate representations for compressing logs. Rodrigues et al. (Rodrigues et al., 2021) propose the lossless compression tool CLP, aiming to quickly retrieve log data while meeting compression requirements. CLP proposes to combine domain-specific compression and search with a generic lightweight compression algorithm. Li et al. (Li et al., 2024d) conduct empirical research on log data and propose LogShrink to overcome their observed limitations by leveraging the commonality and variability of log data. LogBlock (Yao et al., 2021) is designed to help existing jobs perform better. It reduces duplicate logs by preprocessing log headers and rearranging log contents, thereby improving the compression ratio of log files. LogReducer (Yu et al., 2023) is a framework that combines log hotspot identification and online dynamic log filtering. Its non-intrusive design significantly reduces log storage and runtime overhead. $\mu$Slope (Wang et al., 2024) is a compression and search method for semi-structured log data. It achieves efficient storage and query performance through data segmentation, pattern extraction, and index-free design. Denum (Yu et al., 2024) significantly improves log compression rates by optimizing the compression of digital tokens in logs. It is an efficient log compression tool suitable for scenarios where you need to save storage space or transmission bandwidth.
@@ 4.3. Log Parsing
Log data often originates from multiple different devices such as terminals, sensors, and network devices. To analyze it, log parsers are employed to format them into structured and unified ones. Log parsing is usually executed by data classification and template extraction. Data classification is to classify log data into several groups. Each group constitutes a template for extracting features from log data and constructing the structured logs. As shown in Figure 5, the existing log parsers can be taxonomized into 3 categories: clustering-based, pattern-based, and heuristic-based parsers.
@@
Clustering-based parsers classify data using clustering algorithms for log parsing. Xiao et al. (Xiao et al., 2023) propose LPV, which employs a hierarchical clustering algorithm to incrementally group logs based on Euclidean distance. Hamooni et al. (Hamooni et al., 2016) present a rapid log pattern recognition approach named LogMine. It is implemented in the map-reduce framework for distributed platforms to process millions of log messages in seconds. LogCluster (Lin et al., 2016) reduces the number of logs that need to be manually checked and improves the accuracy of problem identification through log clustering and the use of knowledge bases. METING (Coustié et al., 2020) provides a robust and efficient log parsing method through frequent n-gram mining and flexible log grouping strategy, which can effectively process various types of log data.
@@
Frequency-based parsers discover patterns that exceed the frequency threshold and employ the mined patterns to parse logs. Sedki et al. (Sedki et al., 2022) propose the log parsing tool ULP, which combines string matching and local frequency analysis to efficiently parse large log files. Dai et al. (Dai et al., 2020) propose Logram, which utilizes an n-gram dictionary for log parsing. For n-grams with a frequency below the threshold, Logram recursively converts to (n-1)-grams until a list of uncommon 2-grams is obtained. To mitigate the parameter sensitivity issue in log parsers, Dai et al. (Dai et al., 2023) further proposed an entropy-based log parser PILAR, which balances parsing accuracy and efficiency. Xu et al. (Xu et al., 2023a) propose a hybrid log parsing model called Hue, which performs parsing through user-adaptive methods. Prefix-Graph (Chu et al., 2021) is an efficient, adaptive, and universal log parsing method that can stably extract log templates without relying on domain knowledge and manual parameter tuning.
@@
Heuristic-based parsers rely on empirical knowledge to classify log data. He et al. (He et al., 2017) propose the online log parsing method Drain, which employs a depth-fixed parsing tree to group the original logs and encodes them using specially designed parsing rules. Le et al. (Le and Zhang, 2023) propose to use a hint-based few-sample learning algorithm, LogPPT, to capture log template patterns. Utilizing new prompt tuning methods and an adaptive random sampling algorithm, LogPPT performs well on multiple public datasets. Liu et al. (Liu et al., 2022b) propose the UniParser parser to address the issue of difficult processing of heterogeneous logs, using the Token Encoder and Context Encoder modules to learn log context features. Spell (Du and Li, 2016) is an efficient streaming log parsing method that can dynamically extract log patterns in online processing and significantly improve processing efficiency through pre-filtering steps. Logan (Agrawal et al., 2019) achieves efficient and scalable log parsing through distributed processing, LCS matching, dynamic matching tolerance, and periodic merging. USTEP (Vervaet et al., 2021) is an online log parsing method based on an evolutionary tree structure that can discover and encode new parsing rules. It achieves constant parsing time and can efficiently parse raw log messages in a streaming manner.
@@ 5. Intrusion Detection
The intrusion detection stage uncovers intrusions relying on the semantic-level information. This section classifies and summarizes the mainstream graph summarization (Section 5.1), attack detection (Section 5.2), and attack investigation (Section 5.3).
@@ 5.1. Graph Summarization
It is illustrated that stealthy malware will inevitably interact with the underlying OS and be captured by provenance monitoring systems (Wang et al., 2020), which is the reason why PIDS (a form of DL-IDS) has worked and flourished recently. Log data generated from provenance monitoring systems is referred to as data provenance as mentioned. Offering advantages in high precision, data provenance sacrifices memory performance to record all trails of events from their creations to their current states, even some of which are trivial. Unlike network traffic and application logs, data provenance is fine-grained, detailed, and rich in semantics. As a result, the token-level or byte-level log storage systems (Section 4.2.1) and log compression algorithms (Section 4.2.2) are insufficient to handle the memory efficiency of data provenance due to the absence of semantic-level information.
To this end, graph summarization is investigated to further reduce the size of log data semantically. In graph summarization, data provenance is transformed into a provenance graph, of which the causal relations are utilized to build the semantic understanding of system activities. Referring to the definition of data provenance (Definition 3.7), provenance graph is defined as follows:
@@
Provenance graphs allow graph summarization approaches to reduce the size of log data by confidently removing irrelevant events, aggregating similar events, gathering similar execution entities, etc. This categorizes them as a type of lossy reduction, yet the aforementioned log storage and compression are usually lossless (except for statistical-based log compression). We note that some surveys (e.g., (Zipperle et al., 2022; Inam et al., 2023)) may interchangeably use graph summarization and log compression to identify the approaches that reduce the size of log data. In this work, we explicitly distinguish them and refer to the lossless reduction as compression and the opposite one as summarization. Table 3 presents the overview of graph summarization approaches. We classify them into two categories: offline graph summarization and online graph summarization.